Circular No. 7309

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999ew IN NGC 3677 M. Modjaz and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered CCD images taken on Nov. 13.5 and 14.5 UT, both showing magnitude ∼ 16.5. SN 1999ew is located at α = 11^{h}26^{m}17^{s}.66, δ = +46^{o}58′32′′.7 (equinox 2000.0), which is 0′′.3 west and 6′′.1 north of the nucleus of NGC 3677. A KAIT image of the same field on May 23.3 (limiting mag ∼ 19.0) showed nothing at the position of SN 1999ew.

P. Garnavich, R. Kirshner, and S. Jha, Harvard-Smithsonian Center for Astrophysics; and J. Tonry, University of Hawaii (UH), report: “A spectrum (range 550–800 nm) of SN 1999ew, obtained on Nov. 14.6 UT with the Keck II telescope (+ LRIS), shows this to be a type-II supernova, at or before maximum light. The spectrum displays broad Hα and He I 587.5nm features with P-Cyg profiles. The continuum rises steeply into the blue. The photospheric expansion velocity, measured from the minimum of the Hα absorption, is 11 500 km/s. CCD imaging with the UH 2.5-m telescope, obtained simultaneously with the spectroscopy, provides magnitudes V ∼ 17.2, R ∼ 16.9. The redshift of NGC 3677 and the early phase at discovery of SN 1999ew makes it an excellent supernova for application of the Expanding Photosphere Method (Eastman et al. 1996, Ap.J. 466, 911).”

GRS 1915+105 E. B. Waltman, Naval Research Laboratory; and F. D. Ghigo, National

Radio Astronomy Observatory, report: “The Green Bank Interferometer (GBI) has observed a flare in the microquasar GRS 1915+105 on Nov.

14. Flux densities reached 0.4 Jy at 2 GHz on Nov. 15.0 UT. The data are in the public domain and may be accessed via the GBI web page at

http://www.gb.nrao.edu.”

COMET C/1999 S3 (LINEAR) Visual m_{1} estimates: Oct. 29.77 UT, 12.1 (M. Meyer, Frauenstein,

Germany, 0.25-m reflector); Nov. 2.30, 12.4 (N. Biver, Oahu, HI, 0.26-m reflector); 6.58, 11.1 (S. Yoshida, Ibaraki, Japan, 0.25-m reflector); 13.45, 11.5 (Y. Nagai, Yamanashi, Japan, 0.32-m reflector).

1999 November 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green